OpenStudy (anonymous):

supply and demand functions S(x) and D(x), are given for a particular commodity in terms of the level of production x find the value of x[e]. S(x)=5x+280 and D(x)=-5x+480 How can I get the P[e]? Why is = 380. Please help!! Thanks!

OpenStudy (anonymous):

hate economics and dont remember much im assuming you're trying to determine the point when demand is equal to supply curve so d(x)=S(x) and solve for x hint: x=20 once you determine x, input that back into either equation to get your answer

OpenStudy (anonymous):

I am taking an online class and it doesn't show how? Thats why i am lost.

OpenStudy (anonymous):

5x+280 =-5x+480 x=20 d(20)=-5*20+480 =?

OpenStudy (anonymous):

it helps to draw a picture

OpenStudy (anonymous):

ohh waoo, I was trying to do it now.. Thx